OpenStudy (anonymous):

Solve for d. -10 = -2 + 1/2 d

OpenStudy (anonymous):

-10=-2+(1/2)d -8=d/2 -16=d

OpenStudy (anonymous):

-12 + -2 = -10

OpenStudy (anonymous):

-12+-2=-14....